OpenStudy (anonymous):

If the base of a triangle is 8cm greater than the height, and the area is 42 cm^2, what is the height of the triangle and the length of the base?

Directrix (directrix):

A = 1/2 bh where b is the base and h is the height drawn to that base. 42 = 1/2 (h + 8) h 84 = (h + 8) h Next Step?